# GaussianEM: Model compositional and conformational heterogeneity using 3D Gaussians

### v1

@@ Abstract
Understanding protein flexibility and its dynamic interactions with other molecules is essential for protein function study. Cryogenic electron microscopy (cryo-EM) provides an opportunity to directly observe macromolecular dynamics. However, analyzing datasets that contain both continuous motions and discrete states remains highly challenging. Here we present GaussianEM, a Gaussian pseudo-atomic framework that simultaneously models compositional and conformational heterogeneity from experimental cryo-EM images. GaussianEM employs a two-encoder-one-decoder architecture to map an image to its individual Gaussian components, and represent structural variability through changes in Gaussian parameters. This approach provides an intuitive and interpretable description of conformational changes, preserves local structural consistency along the transition trajectories, and naturally bridges the gap between density-based models and corresponding atomic models. We demonstrate the effectiveness of GaussianEM on both simulated and experimental datasets.
Background
@@
Cryogenic electron microscopy (cryo-EM) is a leading technique for imaging biological macromolecules at near-atomic resolution. The structural variability of these molecules is fundamental to their biological functions. In single particle analysis, plunge-frozen images of numerous individual complexes contain rich molecular information, that reflects their compositional and/or conformational heterogeneity lyumkis2019challenges; glaeser2016good; tang2023conformational. However, the high level of experimental noise makes it difficult to extract meaningful structural information from a single snapshot. To address this, averaging is typically combined with classification to divide the cryo-EM dataset into multiple structurally similar subsets, enabling the reconstruction of stable and reliable three-dimensional structures scheres2016processing; scheres2012relion; punjani2017cryosparc; grant2018cis. These approaches have proven effective for datasets exhibiting a discrete number of compositional variation, where each class contains a sufficient number of particles for reconstruction. Multi-body refinement nakane2018characterisation further improves local variation estimation by dividing complexes into independently rigid-moving components ilca2015localized; xu2021conformational; holvec2024structure. However, the limitation on the size of the independent parts restricts its application to highly flexible datasets or motions of small subunits.
@@
Recently, advanced neural networks have been introduced to uncover detailed and complex structural variations in cryo-EM data. CryoDRGN zhong2021cryodrgn, was the first to map projection images to a continuous latent space using a neural radiance field (NeRF) framework. It thereby eliminates the need for user-defined discrete structural classes. Given particle poses obtained from homogeneous reconstruction, CryoDRGN bridges experimental images and reconstructed volumes in the Fourier domain according to Fourier-slice theorem. By traversing the continuous latent space, macromolecular movies can be generated to visualize protein motions and capture both compositional and conformational heterogeneity. Its results demonstrate expressive potential for estimating structural variation from a single cryo-EM dataset. Building upon the NeRF architecture, several subsequent approaches luo2023opus; herreros2025real represent 3D volumes directly in real space to achieve higher-resolution reconstructions. These methods also design different disentanglement strategies for properties such as particle poses and contrast transfer function (CTF) parameters, aiming to obtain a smoother and more physically meaningful latent space. CryoSTAR li2024cryostar further incorporates structural priors and constraints from protein atomic models to enhance reconstruction quality and regulate the variation scale within the estimated conformational landscape. However, NeRF-based methods inherently require explicit 3D volume reconstruction for map comparison and are unable to directly visualize continuous structural transformations. Therefore, interpreting multiple points across the estimated conformation landscape is often time consuming and relies heavily on manual inspection. In particular, local structures often lack consistency within the reconstructed serial volumes.
@@
e2gmm chen2021deep in EMAN2, represents the protein structure using a Gaussian mixture model (GMM) and intuitively describes structural heterogeneity through variations in Gaussian parameters. It measures the difference between experimental images and the GMM projection in terms of Fourier Ring Correlation (FRC) values, and constructs network inputs by stacking the FRC gradients with respect to the GMM. The explicit modeling of the existence and displacement of Gaussian spheres facilitates the localization of structural changes. However, the use of stacked gradient inputs constrains the number of Gaussians that can be handled for high-resolution reconstruction and reduces the model’s sensitivity to complex motions. DynaMight schwab2023dynamight focuses on continuous motion estimation by optimizing Gaussian displacements in the Fourier domain. It additionally estimates inverse deformations to improve 3D consensus reconstruction and supports a larger number of Gaussian spheres derived from atomic models. 3DFlex punjani20233dflex directly estimates a convection field to deform the 3D consensus structure, enabling the simultaneous generation of high-resolution 3D volumes and deformation fields corresponding to experimental images. However, both methods are limited in their ability to characterize compositional heterogeneity. Consequently, intuitive interpretation of both discrete and continuous conformational changes remains a challenge.
@@
To evaluate its performance, we conducted comprehensive experiments on multiple public datasets. The results reveal that GaussianEM effectively and intuitively captures both discrete and continuous conformational changes from experimental images. Additionally, it preserves the integrity and consistency of local structures during continuous transitions along the conformational trajectory.
@@ Results
In GaussianEM, we assume that the heterogeneity present in experimental particle images can be represented by a single consensus structure along with its corresponding structural variations. Compositional heterogeneity is manifested as partial structural missing, whereas conformational heterogeneity appears as structural displacement. Figure 1a illustrates the architecture of GaussianEM. The method uses a set of 3D Gaussians to fit the consensus model and identifies parameter variations of Gaussians by learning an adapted variation autoencoder (VAE). Each Gaussian is defined by seven parameters: a density value, a 3D scale vector used to compute its covariance, and 3D spatial coordinates. Local structural motions are described by changes in the 3D positions of Gaussians that make up the model, and the presence or absence of components is reflected by variations in their density values. The initial Gaussians are uniformly sampled from the homogeneous reconstruction, filtered by provided contour level. The number of Gaussians determines the achievable resolution of reconstructed structures, but also has a substantial impact on computational cost. In practice, GaussianEM adopts a sampling interval of two voxels, which works well on a wide range of experimental datasets we have used.
@@
The heterogeneous reconstruction network consists of two encoders, an image encoder and a Gaussian encoder, and a single decoder that integrates their representations. The image encoder processes 2D particle images and transforms them into a continuous low-dimensional latent conformation space. Each point embedded into this low-dimension manifold corresponds to a potential 3D conformation. Although the number of Gaussians is considerably smaller than that of voxels, directly inputting all Gaussian parameters into the network remains computationally infeasible. To address this challenge, GaussianEM introduces a learnable Gaussian encoding submodule to allow the network to capture the contextual relationships among Gaussians. The Gaussian encoder accepts attributes such as density, scale, and position for each Gaussian, and outputs a unique latent embedding that serves as its identifier. Inspired by NeRF-based architectures, positional encoding is incorporated to help the network capture high-frequency spatial information, thereby enhancing its ability to reconstruct fine structural details. The latent variable from the image encoder is then concatenated with the Gaussian embedding and passed to the decoder, which predicts the parameter variations for each Gaussian. This per-Gaussian paradigm allows efficient batch-level parallelization and facilitates the aggregation and discrimination of Gaussians according to their learned parameter changes.
@@
In GaussianEM, the difference between the experimental particle images and the rendered projections (given known particle poses) is measured in the real-space domain. Although we still need to transform the projection twice to apply the CTF modulator, comparison in real space makes the resulting structural changes far more intuitive and interpretable. Once the network is trained, the learned conformation space can be visualized using standard dimensionality reduction techniques mcinnes2018umap. By identifying cluster centers within this continuous latent space, representative 3D volumes can be generated to capture discrete conformational states. Moreover, since the conformation manifold is continuous, trajectories between clusters can be smoothly traversed to produce dynamic structural transitions. An additional advantage of the real-space Gaussian representation is its natural compatibility with atomic models. The displacements or loss of Gaussians are easily mapped to the atomic model through nearest point connection. This property facilitates the generation of deformed atomic models (PDB structures) that correspond to different conformational states inferred from the cryo-EM data. Consequently, GaussianEM provides an optional perspective that links density-based heterogeneity analysis with atomic-level structures.
@@
We applied GaussianEM to one simulation CryoBench dataset jeon2024cryobench and six publicly available Electron Microscopy Public Image Archive (EMPIAR) iudin2016empiar datasets. Four experimental datasets are presented in the main text, and two additional datasets are included in the Supplementary Material (sections $\alpha V\beta 8$ integrin bound to latent TGF-$\beta$” and “TRPV1 ion channel”). For each dataset, consensus reconstruction and nonuniform refinement were performed in CryoSPARC punjani2017cryosparc using all available experimental images. The estimated particle poses were fixed during training, and particle images were downsampled to a uniform size of $256\times 256$. For datasets with smaller raw image sizes, the original resolution was retained. Initial Gaussian parameters were sampled from the consensus map after applying a contour-level threshold. When a corresponding PDB atomic model was available, Gaussian displacements were transferred to atomic coordinates to generate conformation-specific atomic models. The reconstructed density maps and atomic models were visualized in ChimeraX pettersen2021ucsf. All experiments were conducted on an NVIDIA GeForce RTX 3080 with 10 GB RAM, with training typically completed within 3 to 10 hours, depending on the scale of the dataset.
@@
Simulation data of large scale motion
@@
The simulated dataset of the human immunoglobulin G (IgG) antibody complex from CryoBench was used to evaluate the ability of GaussianEM to capture large-scale molecular motions. This dataset models a one-dimensional continuous circular motion by rotating the dihedral angle that connects one of the fragment antibody (Fab) domains to the rest of the molecule. It includes 100 conformations, each with corresponding density maps and atomic models, and a total of 100,000 particle images with a box size of 128 pixels of width 3.0 Å. Gaussian-based method e2gmm failed to describe this large-scale motion (see Figure S1 in Supplementary Material section “IgG-1D simulation”). In contrast, GaussianEM successfully recovered the complete 360 rotational motion, as illustrated by the reconstructed volumes in Figure 2b. The latent conformation space exhibits a distinct circular distribution, consistent with the continuous 3D rotational motion of the molecule. The motion is shown from two different viewing angles, including a top view aligned with the rotation axis. The reconstructed maps maintain comparable resolution across the full range of large-scale motions. Four representative conformations are displayed, each rotated by 90 along the rotational axis.
@@
We further mapped the estimated Gaussian displacements onto the atoms associated with the consensus structure, as illustrated by the red dashed lines in Figure 2c. Unlike multi-body refinement, GaussianEM does not explicitly impose strict rigid-body constraints on the deformation fields, but the resulting atomic displacements still exhibit physically consistent motion. The Gaussian encoding paradigm, together with the local rigidity regularization term, collectively encourages neighboring Gaussians to undergo similar variations. While voxel-based methods can also capture rotational motions, they sometimes introduce inconsistencies in local regions (see Figure S1 in the Supplementary Material, section “IgG-1D simulation”). In contrast, GaussianEM preserves the structural continuity of deformed volumes, as all deformations are applied on the same consensus model. The two superimposed atomic models show a high degree of structural overlap, with a RMSD of 2.26 Å  between corresponding atoms. This result indicates that GaussianEM can accurately capture large-scale conformational motion. Moreover, we reconstructed 1000 atomic models to measure the RMSD distribution, as shown in Figure 2d. When categorized by rotation angles, the results show that conformations closer to the original structure yield smaller RMSD values, indicating higher reconstruction accuracy. For composite conformations within the 144–180 range, the average RMSD is approximately 8 Å, which remains sufficient for capturing conformational heterogeneity.
@@
The bL17-limited E. coli ribosome assembly intermediates (EMPIAR-10841) rabuck2022quantitative contains a large number of discrete conformations, which have been validated by both classification-based and deep learning–based analyses zhong2021cryodrgn; rabuck2022quantitative. The dataset consists of 123804 particle images with a box size of 160 pixels of width 2.64 Å. In our experiment, we first allowed both the density and position parameters of the Gaussians to vary, without applying any masking operation. Uniform Manifold Approximation and Projection (UMAP) mcinnes2018umap was used to reduce the latent variables to 2D space to visualize the conformation distributions. The resulting latent space shows a clear separation of particle clusters, from which we select ten representative cluster centers for detailed analysis. As shown in Figure 3b, distinct structural differences are observed among the clusters, while subtle compositional changes are highlighted by the red boxes. The large-scale structural differences likely involve the ribosomal base and several stalk regions, reflecting different stages or pathways of ribosome assembly rabuck2022quantitative. The presence of smaller local regions may correspond to the incorporation of ribosomal RNA (rRNA) segments during the maturation process. Class 9 appears distant from other clusters in the 2D latent space, and its reconstructed volume exhibits a different orientation, likely due to minor errors in pose estimation. We further investigated subtle structural variability within the dataset, as illustrated in Figure 3c. The example highlights ribosomal proteins L16 and L25, which appear in varying states—fully present, partially present, or completely absent—across different conformations.
Following the smearing effect observed in the dynamic central protuberance (CP) domain chen2021deep, we subsequently constrained the variation to position parameters within that region. A clear tilting motion of that domain is revealed, as shown in Figure 3d. For better visualization, we aligned the PDB 6PJ6 atomic model to the reconstructed volume and map the Gaussian displacements onto the corresponding atoms. The superimposed atomic models are displayed alongside the reconstructed volumes, illustrating the detailed conformational changes captured by GaussianEM. We further analyzed the learned Gaussian embeddings by performing k-means clustering. Four clusters were identified and colored in Figure 3e. By removing each cluster in turn, we observed that the remaining structure closely resembled the corresponding discrete conformations, indicating that the Gaussian embeddings effectively capture different structural features.
@@
GaussianEM was further applied to analyze structural variations in the pre-catalytic spliceosome dataset (EMPIAR-10180) plaschka2017structure. This dataset has 327490 particle images with a box size of 320 pixels of width 1.699 Å. The images were downsampled to $256\times 256$, and initial Gaussians were sampled from the consensus map deposited in EMPIAR-10180. All Gaussian parameters were allowed to vary during training. Using the K-means algorithm, six distinct clusters were identified (Figure 4a). Reconstruction of density maps from the cluster centers revealed clear compositional differences and continuous motion trajectories. Notably, the absence of the SF3b and foot domains is highlighted by the red boxes in Figure 4b. Similar compositional heterogeneity was captured by the NeRF-based approach OPUS-DSD luo2023opus but not by the Gaussian-based method e2gmm chen2021deep. Clustered projections were further reconstructed in CryoSPARC to validate the effectiveness of GaussianEM (Figure 4c). The absence of SF3B and foot domains is also clearly visible. The learned Gaussian embeddings were grouped into five clusters, as shown in Figure 4d. Each cluster roughly corresponds to a distinct biological domain. The blue cluster predominantly represents the SF3b region, while the purple cluster mainly corresponds to the helicase domain. Furthermore, PCA of the latent space reveals distinct patterns of conformational variability along the first two principal components (PC1–PC2) in Figure 4e. The helicase and SF3b domains exhibit coordinated but distinct rotational motions along different components. Along the first principal component, the spliceosome transitions from an “open” to a “closed” state: at the beginning of the trajectory, the SF3b domain stays on the top of the core body, and at the end, it folds downward toward the core region. Along the second principal component, the helicase and SF3b domains jointly display a left–right swinging motion. The displacements of Gaussians were mapped to the atomic model PDB-5NRL, which was fitted to density maps by CryoAlign he2024accurate. The resulting atomic motions were smooth and coherent, indicating that our method can suppress artificial structural distortions to some extent. Voxel-based methods can produce overall continuous trajectories but may exhibit local inconsistencies (see Figure S2 in Supplementary Material section “Pre-catalytic spliceosome”).
@@
Figure 5: Heterogeneity analysis of BtpeA-BtaeB-VgrG complex. a Generation of the initial Gaussians. Valid points were extracted from two density maps (EMD-61469 and EMD-61458) and combined to form the initial model. b Three reconstructed volumes along the first principal component. A region not present in publicly available data is highlighted by a red box. c Reconstructed maps by CryoSPARC using clustered projection from GaussianEM. d UMAP visualization of principal components 2 and 3. The reconstructed volumes and their corresponding atomic models are shown on the right.
@@
We used the cargo complex (BtpeA-BtaeB-BtapC) bound to the VgrG spike li2025conserved as the fourth test. This dataset consists of 106565 particle images with a box size of 480 pixels of width 1.18 Å. In contrast to the previous datasets, this dataset lacked a complete density for Gaussian initialization. Two related maps, EMD-61458 and EMD-61459, each represented partial structures. The homogeneous reconstruction from CryoSPARC also provided additional structural information. We extracted valid points from these maps and aligned them using CryoAlign to generate a unified initial Gaussian model. The particle images were downsampled to $256\times 256$, and density and position parameters of Gaussians were permitted to change.

### v2

# Resolving compositional and conformational heterogeneity in cryo-EM with deformable 3D Gaussian representations

@@ Abstract
Understanding protein flexibility and its dynamic interactions with other molecules is essential for studying protein function. Although cryogenic electron microscopy (cryo-EM) provides an opportunity to observe macromolecular dynamics directly, computational analysis of datasets mixing continuous and discrete structural states remains a formidable challenge. Here we introduce GaussianEM, a Gaussian-based pseudo‑atomic framework that simultaneously resolves compositional and conformational heterogeneity from cryo‑EM images. GaussianEM employs a dual-encoder–single-decoder architecture to decompose images into learnable Gaussian components, with variability encoded through modulated parameters. This explicit parameterization yields a continuous, intuitive representation of conformational dynamics that inherently preserves local structural integrity. By modeling displacements in Gaussian space, we capture atomic‑scale conformational landscapes, bridging density maps and all‑atom models. In comprehensive experiments, GaussianEM successfully reconstructs complex compositional and conformational variability, and resolves previously unobserved details in public datasets. Quantitative evaluations further confirm its ability to capture broader conformational diversity without sacrificing structural fidelity.
Introduction
@@
Cryogenic electron microscopy (cryo-EM) is a leading technique for imaging biological macromolecules at near-atomic resolution. The structural variability of these molecules is fundamental to their biological functions. In single particle analysis, plunge-frozen images of numerous individual complexes contain rich molecular information that reflects their compositional and/or conformational heterogeneity Lyumkis (2019); Glaeser (2016); Tang et al. (2023). However, the high level of experimental noise makes it difficult to extract meaningful structural information from a single snapshot. To address this, averaging is typically combined with classification to divide the cryo-EM dataset into multiple structurally similar subsets, enabling the reconstruction of stable and reliable three-dimensional structures Scheres (2016, 2012); Punjani et al. (2017); Grant et al. (2018). These approaches have proven effective for datasets exhibiting a discrete number of compositional variations, where each class contains a sufficient number of particles for reconstruction. Multi-body refinement Nakane et al. (2018) further improves local variation estimation by dividing complexes into independently rigid-moving components Ilca et al. (2015); Xu et al. (2021); Holvec et al. (2024). However, the limitation on the size of the independent parts restricts its application to highly flexible datasets or motions of small subunits.
@@
Recently, advanced neural networks have been introduced to uncover detailed and complex structural variations in cryo-EM data. CryoDRGN Zhong et al. (2021) was the first to map projection images to a continuous latent space using a neural radiance field (NeRF) framework. It thereby eliminates the need for user-defined discrete structural classes. Given particle poses obtained from homogeneous reconstruction, CryoDRGN bridges experimental images and reconstructed volumes in the Fourier domain according to the Fourier-slice theorem. By traversing the continuous latent space, macromolecular movies can be generated to visualize protein motions and capture both compositional and conformational heterogeneity. Its results demonstrate expressive potential for estimating structural variation from a single cryo-EM dataset. Building upon the NeRF architecture, several subsequent approaches Luo et al. (2023); Herreros et al. (2025) represent 3D volumes directly in real space to achieve higher-resolution reconstructions. These methods also design different disentanglement strategies for properties such as particle poses and contrast transfer function (CTF) parameters, aiming to obtain a smoother and more physically meaningful latent space. CryoSTAR Li et al. (2024) further incorporates structural priors and constraints from protein atomic models to enhance reconstruction quality and regulate the variation scale within the estimated conformational landscape. However, NeRF-based methods inherently require explicit 3D volume reconstruction for map comparison and are unable to directly visualize continuous structural transformations. Therefore, interpreting multiple points across the estimated conformation landscape is often time-consuming and relies heavily on manual inspection. In particular, local structures often lack consistency within the reconstructed serial volumes.
@@
e2gmm Chen and Ludtke (2021) in EMAN2, represents the protein structure using a Gaussian mixture model (GMM) and intuitively describes structural heterogeneity through variations in Gaussian parameters. It measures the difference between experimental images and the GMM projection in terms of Fourier Ring Correlation (FRC) values, and constructs network inputs by stacking the FRC gradients with respect to the GMM. The explicit modeling of the existence and displacement of Gaussian spheres facilitates the localization of structural changes. However, the use of stacked gradient inputs constrains the number of Gaussians that can be handled for high-resolution reconstruction and reduces the model’s sensitivity to complex motions. DynaMight Schwab et al. (2023) focuses on continuous motion estimation by optimizing Gaussian displacements in the Fourier domain. It additionally estimates inverse deformations to improve 3D consensus reconstruction and supports a larger number of Gaussian spheres derived from atomic models. 3DFlex Punjani and Fleet (2023) directly estimates a convection field to deform the 3D consensus structure, enabling the simultaneous generation of high-resolution 3D volumes and deformation fields corresponding to experimental images. However, both methods are limited in their ability to characterize compositional heterogeneity. Consequently, intuitive interpretation of both discrete and continuous conformational changes remains a challenge.
@@
To evaluate the performance of GaussianEM, we conducted comprehensive experiments on multiple simulated benchmarks and EMPIAR datasets. The results demonstrate that GaussianEM effectively and intuitively captures both discrete and continuous conformational changes in the ribosome assembly, the pre-catalytic spliceosome, the $\alpha V\beta 8$ integrin, and the $\alpha V\beta 8$ integrin bound to latent TGF-$\beta$. Notably, GaussianEM identifies novel conformational states of T6SS effectors that were not resolved in previously published analyses. Additionally, GaussianEM preserves the integrity and consistency of local structures during continuous transitions along the conformational trajectory, as it explicitly models deformations applied to the consensus Gaussian model. Quantitative evaluations focusing on conformational diversity demonstrate GaussianEM’s distinct advantage over CryoDRGN, DynaMight, and 3DFlex in capturing a wider range of conformational states. Furthermore, FSC-based comparisons on benchmark datasets confirm the high fidelity and reliability of GaussianEM’s reconstruction results.
@@ Results
In GaussianEM, we assume that the heterogeneity present in experimental particle images can be represented by a single consensus structure along with its corresponding structural variations. Compositional heterogeneity is manifested as partial structural missing, whereas conformational heterogeneity appears as structural displacement. Figure 1a illustrates the architecture of GaussianEM. The method uses a set of 3D Gaussians to fit the consensus model and identifies parameter variations of Gaussians by learning an adapted variation autoencoder (VAE). Each Gaussian is defined by seven parameters: a density value, a 3D scale vector used to compute its covariance, and 3D spatial coordinates. Local structural motions are described by changes in the 3D positions of Gaussians that make up the model, and the presence or absence of components is reflected by variations in their density values. The initial Gaussians are uniformly sampled from the homogeneous reconstruction, filtered by the provided contour level. The number of Gaussians determines the achievable resolution of reconstructed structures, but also has a substantial impact on computational cost. In practice, GaussianEM adopts a sampling interval of two voxels, which works well on a wide range of experimental datasets we have used.
@@
The heterogeneous reconstruction network consists of two encoders, an image encoder and a Gaussian encoder, and a single decoder that integrates their representations. The image encoder processes 2D particle images and transforms them into a continuous low-dimensional latent conformation space. Each point embedded into this low-dimensional manifold corresponds to a potential 3D conformation. Although the number of Gaussians is considerably smaller than that of voxels, directly inputting all Gaussian parameters into the network remains computationally infeasible. To address this challenge, GaussianEM introduces a learnable Gaussian encoding submodule to allow the network to capture the contextual relationships among Gaussians. The Gaussian encoder accepts attributes such as density, scale, and position for each Gaussian, and outputs a unique latent embedding that serves as its identifier. Inspired by NeRF-based architectures, positional encoding is incorporated to help the network capture high-frequency spatial information, thereby enhancing its ability to reconstruct fine structural details. The latent variable from the image encoder is then concatenated with the Gaussian embedding and passed to the decoder, which predicts the parameter variations for each Gaussian. This per-Gaussian paradigm allows efficient batch-level parallelization and facilitates the aggregation and discrimination of Gaussians according to their learned parameter changes.
@@
In GaussianEM, the difference between the experimental particle images and the rendered projections (given known particle poses) is measured in the real-space domain. Although we still need to transform the projection twice to apply the CTF modulator, comparison in real space makes the resulting structural changes far more intuitive and interpretable. Once the network is trained, the learned conformation space can be visualized using standard dimensionality reduction techniques McInnes et al. (2018). By identifying cluster centers within this continuous latent space, representative 3D volumes can be generated to capture discrete conformational states. Moreover, since the conformation manifold is continuous, trajectories between clusters can be smoothly traversed to produce dynamic structural transitions. An additional advantage of the real-space Gaussian representation is its natural compatibility with atomic models. The displacements or loss of Gaussians are easily mapped to the atomic model through the nearest point connection. This property facilitates the generation of deformed atomic models (PDB structures) that correspond to different conformational states inferred from the cryo-EM data. Consequently, GaussianEM provides an optional perspective that links density-based heterogeneity analysis with atomic-level structures.
@@
We applied GaussianEM to two simulation CryoBench datasets Jeon et al. (2024) and five publicly available Electron Microscopy Public Image Archive (EMPIAR) Iudin et al. (2016) datasets. Four experimental datasets are presented in the main text, and an additional datasets are included in the Supplementary Material (sections $\alpha V\beta 8$ integrin bound to latent TGF-$\beta$”). For each dataset, consensus reconstruction and nonuniform refinement were performed in CryoSPARC Punjani et al. (2017) using all available experimental images. The estimated particle poses were fixed during training, and particle images were downsampled to a uniform size of $256\times 256$. For datasets with smaller raw image sizes, the original resolution was retained. Initial Gaussian parameters were sampled from the consensus map after applying a contour-level threshold. When a corresponding PDB atomic model was available, Gaussian displacements were transferred to atomic coordinates to generate conformation-specific atomic models. The reconstructed density maps and atomic models were visualized in ChimeraX Pettersen et al. (2021). All experiments were conducted on an NVIDIA GeForce RTX 3080 with 10 GB RAM, with training typically completed within 3 to 10 hours, depending on the scale of the dataset.
@@
Simulation data of large-scale motion
@@
The simulated IgG-1D dataset of the human immunoglobulin G (IgG) antibody complex from CryoBench was used to evaluate the ability of GaussianEM to capture large-scale molecular motions. This dataset models a one-dimensional continuous circular motion by rotating the dihedral angle that connects one of the fragment antibody (Fab) domains to the rest of the molecule. It includes 100 conformations, each with corresponding density maps and atomic models, and a total of 100,000 particle images with a box size of 128 pixels of width 3.0 Å. Gaussian-based method e2gmm failed to describe this large-scale motion (see Figure LABEL:fig:IgG in Supplementary Material section “IgG simulation”). In contrast, GaussianEM successfully recovered the complete 360 rotational motion, as illustrated by the reconstructed volumes in Figure 2b. The latent conformation space exhibits a distinct circular distribution, consistent with the continuous 3D rotational motion of the molecule. The motion is shown from two different viewing angles, including a top view aligned with the rotation axis. The reconstructed maps maintain comparable resolution across the full range of large-scale motions. Four representative conformations are displayed, each rotated by 90 along the rotational axis.
@@
We further mapped the estimated Gaussian displacements onto the atoms associated with the consensus structure, as illustrated by the red dashed lines in Figure 2c. Unlike multi-body refinement, GaussianEM does not explicitly impose strict rigid-body constraints on the deformation fields, but the resulting atomic displacements still exhibit physically consistent motion. The Gaussian encoding paradigm, together with the local rigidity regularization term, collectively encourages neighboring Gaussians to undergo similar variations. While voxel-based methods can also capture rotational motions, they sometimes introduce inconsistencies in local regions (see Figure LABEL:fig:IgG in the Supplementary Material section “IgG simulation”). In contrast, GaussianEM preserves the structural continuity of deformed volumes, as all deformations are applied to the same consensus model. The two superimposed atomic models show a high degree of structural overlap, with an RMSD of 2.26 Å  between corresponding atoms. This result indicates that GaussianEM can accurately capture large-scale conformational motion. Moreover, we reconstructed 1000 atomic models to measure the RMSD distribution, as shown in Figure 2d. When categorized by rotation angles, the results show that conformations closer to the original structure yield smaller RMSD values, indicating higher reconstruction accuracy. For composite conformations within the 144–180 range, the average RMSD is approximately 8 Å, which remains sufficient for capturing conformational heterogeneity.
To model more realistic structural heterogeneity, CryoBench also introduces the IgG-RL dataset, which is generated by modeling the 5-residue linker through backbone dihedral angle sampling based on disordered peptide Ramachandran distributions. This dataset comprises 100 conformations, with 1,000 projections generated for each conformation at a box size of 128 pixels and pixel size of 3.0 Å. Voxel-based CryoDRGN displayed occasional loss of density in structural domains of the IgG-RL dataset exhibiting disordered dynamics (see Figure LABEL:fig:IgG_RLc in the Supplementary Material section “IgG simulation”). In contrast, GaussianEM can recover conformational motions exhibiting stochastic variations (see Figure LABEL:fig:IgG_RLb in the Supplementary Material, section “IgG simulation”).
@@
The bL17-limited E. coli ribosome assembly intermediates (EMPIAR-10841) Rabuck-Gibbons et al. (2022) contains a large number of discrete conformations, which have been validated by both classification-based and deep learning–based analyses Zhong et al. (2021); Rabuck-Gibbons et al. (2022). The dataset consists of 123804 particle images with a box size of 160 pixels of width 2.64 Å. In our experiment, we first allowed both the density and position parameters of the Gaussians to vary, without applying any masking operation. Uniform Manifold Approximation and Projection (UMAP) McInnes et al. (2018) was used to reduce the latent variables to the 2D space to visualize the conformation distributions. The resulting latent space shows a clear separation of particle clusters, from which we select ten representative cluster centers for detailed analysis. As shown in Figure 3b, distinct structural differences are observed among the clusters, while subtle compositional changes are highlighted by the red boxes. The large-scale structural differences likely involve the ribosomal base and several stalk regions, reflecting different stages or pathways of ribosome assembly Rabuck-Gibbons et al. (2022). The presence of smaller local regions may correspond to the incorporation of ribosomal RNA (rRNA) segments during the maturation process. Class 9 appears distant from other clusters in the 2D latent space, and its reconstructed volume exhibits a different orientation, likely due to minor errors in pose estimation. We further investigated subtle structural variability within the dataset, as illustrated in Figure 3c. The example highlights ribosomal proteins L16 and L25, which appear in varying states—fully present, partially present, or completely absent—across different conformations.
Following the smearing effect observed in the dynamic central protuberance (CP) domain Chen and Ludtke (2021), we subsequently constrained the variation to position parameters within that region. A clear tilting motion of that domain is revealed, as shown in Figure 3d. For better visualization, we aligned the PDB 6PJ6 atomic model to the reconstructed volume and mapped the Gaussian displacements onto the corresponding atoms. The superimposed atomic models are displayed alongside the reconstructed volumes, illustrating the detailed conformational changes captured by GaussianEM. We further analyzed the learned Gaussian embeddings by performing k-means clustering. Four clusters were identified and colored in Figure 3e. By removing each cluster in turn, we observed that the remaining structure closely resembled the corresponding discrete conformations, indicating that the Gaussian embeddings effectively capture different structural features.
@@
GaussianEM was further applied to analyze structural variations in the pre-catalytic spliceosome dataset (EMPIAR-10180) Plaschka et al. (2017). This dataset has 327490 particle images with a box size of 320 pixels of width 1.699 Å. The images were downsampled to $256\times 256$, and initial Gaussians were sampled from the consensus map deposited in EMPIAR-10180. All Gaussian parameters were allowed to vary during training. Using the K-means algorithm, six distinct clusters were identified (Figure 4a). Reconstruction of density maps from the cluster centers revealed clear compositional differences and continuous motion trajectories. Notably, the absence of the SF3b and foot domains is highlighted by the red boxes in Figure 4b. Similar compositional heterogeneity was captured by the NeRF-based approach OPUS-DSD Luo et al. (2023) but not by the Gaussian-based method e2gmm Chen and Ludtke (2021). Clustered projections were further reconstructed in CryoSPARC to validate the effectiveness of GaussianEM (Figure 4c). The absence of SF3B and foot domains is also clearly visible. The learned Gaussian embeddings were grouped into five clusters, as shown in Figure 4d. Each cluster roughly corresponds to a distinct biological domain. The blue cluster predominantly represents the SF3b region, while the purple cluster mainly corresponds to the helicase domain. Furthermore, PCA of the latent space reveals distinct patterns of conformational variability along the first two principal components (PC1–PC2) in Figure 4e. The helicase and SF3b domains exhibit coordinated but distinct rotational motions along different components. Along the first principal component, the spliceosome transitions from an “open” to a “closed” state: at the beginning of the trajectory, the SF3b domain stays on the top of the core body, and at the end, it folds downward toward the core region. Along the second principal component, the helicase and SF3b domains jointly display a left–right swinging motion. The displacements of Gaussians were mapped to the atomic model PDB-5NRL, which was fitted to density maps by CryoAlign He et al. (2024). The resulting atomic motions were smooth and coherent, indicating that our method can suppress artificial structural distortions to some extent. Voxel-based methods can produce overall continuous trajectories but may exhibit local inconsistencies (see Figure LABEL:fig:10180_drgn in Supplementary Material section “Pre-catalytic spliceosome”).
@@
Figure 5: Heterogeneity analysis of BtpeA-BtaeB-VgrG complex. a Generation of the initial Gaussians. Valid points were extracted from two density maps (EMD-61459 and EMD-61458) and combined to form the initial model. b Three reconstructed volumes along the first principal component. A region not present in publicly available data is highlighted by a red box. c Reconstructed maps by CryoSPARC using clustered projection from GaussianEM. d UMAP visualization of principal components 2 and 3. The reconstructed volumes and their corresponding atomic models are shown on the right.
@@
We then used the cargo complex (BtpeA-BtaeB-BtapC) bound to the VgrG spike Li et al. (2025) to test GaussianEM. This dataset consists of 106565 particle images with a box size of 480 pixels of width 1.18 Å. In contrast to the previous datasets, this dataset lacked a complete density for Gaussian initialization. Two related maps, EMD-61458 and EMD-61459, each represented partial structures. The homogeneous reconstruction from CryoSPARC also provided additional structural information. We extracted valid points from these maps and aligned them using CryoAlign to generate a unified initial Gaussian model. The particle images were downsampled to $256\times 256$, and the density and position parameters of Gaussians were permitted to change.